STARK EFFECT IN X_{2}Y_{4} MOLECULES : APPLICATION TO ETHYLENE

\underline{M. ROTGER}, W. RABALLAND, V. BOUDON, M. LOETE, Laboratoire de Physique de l’Université de Bourgogne, UMR CNRS 5027, 9, av. Alain Savary, B.P. 47870, F-21078 Dijon Cedex, France; J. BREIDUNG, W. THIEL, Max-Planck-Institut für Kohlenforschung, Kaiser-Wilhelm-Platz 1, D-45470 Mülheim an der Ruhr, Germany.

We present a development of the dipole moment and polarizability operators of X_{2}Y_{4} molecules, using a tensorial formalism^{a}^{,}^{b} analogous to the one developed for tetrahedral and octahedral molecules^{c}. These operators are involved in the calculation of the intensities of rovibrational transitions as well as in the calculation of the Stark effect. Expressions for the matrix elements are derived. A model for the study of the Stark effect in isolated bands of such molecules is proposed and has been used to predict the Stark spectra of the ν_{12} band of ethylene. Values of the polarizability coefficients have been calculated using ab initio methods.

aW. Raballand, M. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc. 217, 239-248, (2003)

bCh. Wenger, W. Raballand, M. Rotger and V. Boudon, J. Quant. Spectrosc. Radiat. Transfer, accepted, (2005).

cV. Boudon, J.-P. Champion, T. Gabard, M. Loëte, F. Michelot, G. Pierre, M. Rotger, Ch. Wenger and M. Rey, J. Mol. Spectrosc.

228, 620-634, (2004).